# Optimizing Product Provenance Verification using Data Valuation Methods

### v2

@@ Abstract.
Determining and verifying product provenance remains a critical challenge in global supply chains, particularly as geopolitical conflicts and shifting borders create new incentives for misrepresentation of commodities, such as hiding the origin of illegally harvested timber or agriculture grown on illegally cleared land. Stable Isotope Ratio Analysis (SIRA), combined with Gaussian process regression-based isoscapes, has emerged as a powerful tool for geographic origin verification. However, the effectiveness of these models is often constrained by data scarcity and suboptimal dataset selection. In this work, we introduce a novel data valuation framework designed to enhance the selection and utilization of training data for machine learning models applied in SIRA. By prioritizing high-informative samples, our approach improves model robustness and predictive accuracy across diverse datasets and geographies. We validate our methodology with extensive experiments, demonstrating its potential to significantly enhance provenance verification, mitigate fraudulent trade practices, and strengthen regulatory enforcement of global supply chains.
Data valuation, Data Shapley, Stable Isotope Ratio Analysis (SIRA), Gaussian processes
copyright: none
1. Introduction
@@
Global natural resource supply chains are opaque, especially since natural resources are often transformed from raw materials into finished consumer-facing products. These complex supply chains often involve multiple countries, with intermediate outputs being traded internationally and being used as inputs into further manufacturing processes. In addition to business-commerce decisions driving supply chain sourcing, the economics of natural resource trade are often closely linked with geopolitics. Determining and verifying product provenance is a challenge in global supply chains, as geopolitics and the lure of “don’t ask, don’t tell” with respect to the ecological and social cost creates incentives for misrepresentation of commodities, such as hiding the origin of illegally harvested timber or agriculture grown on illegally cleared land.
Product identification and provenance verification of traded natural resources have emerged as promising research areas, with various combinations of methods used based on the specific natural resource sector and the level of granularity of species identification and origin-provenance determination. For example, for wood and forest products, determining species identification and geographic harvest provenance requires utilizing multiple testing methods and tools (Corruption, 2021; Schmitz et al., 2020; Dormontt et al., 2015). Specific to geographic origin verification, Stable Isotope Ratio Analysis (SIRA), combined with Gaussian process regression-based isoscapes has emerged as a powerful tool (Truszkowski et al., 2025; Mortier et al., 2024).
SIRA can be useful for understanding the origin of organic products. Stable isotopes are chemical variants of elements that do not undergo radioactive decay. The ratio of stable isotopes denotes the relative enrichment of different elemental stable isotopes in a sample which is typically measured by mass spectrometry (Barrie and Prosser, 1996) and allows us to understand the enrichment of these isotopes in that sample. The natural variation observed for this ratio is determined by underlying mechanisms that are affected by a range of different factors including but not limited to environmental, atmospheric, soil, metabolic fraction, or other characteristics specific to a species (Siegwolf et al., 2022; Wang et al., 2021; Vystavna et al., 2021). Previous work has shown how it can be used to trace the origin of items such as timber, seafood, agricultural products, and cotton (Truszkowski et al., 2025; Mortier et al., 2024; Watkinson et al., 2022; Cusa et al., 2022; Wang et al., 2020; Meier-Augenstein et al., 2014).
However, the effectiveness of these models is often constrained by data scarcity and suboptimal dataset selection. The use of SIRA to identify origin is limited by several factors, though chief among them is the limited number of reference samples used as input data for modeling isoscapes. Reference sampling campaigns are expensive, and often sampling locations are chosen based on ease of sampling rather than areas that yield a gain in model prediction accuracy (Gasson et al., 2021). There is a need to optimize reference sample collection efforts, in such a way that would enable the ability to “value” reference samples thus enabling more judicious sample collection campaigns.
Here, we introduce a novel data valuation framework designed to enhance the selection and utilization of training data for machine learning models applied to SIRA. Our key contributions are:
  1. (1)
@@
    We bring the growing literature on data valuation in machine learning to bear upon the pressing problem of optimizing product provenance verification. By prioritizing highly informative samples, our approach improves model robustness and predictive accuracy across diverse datasets and geographies.
  2. (2)
@@
    We have deployed our approach of optimizing the selection of training samples which improves our model accuracy. These new models are being used by European enforcement agencies to stop the trade in sanctioned Russian timber by demonstrating that a claimed harvest location other than Russia is not viable. See coverage of our work from the New York Times  (Nazaryan, 2024). Due to confidentiality reasons, we use a global dataset of Oak (Quercus spp.) reference samples to illustrate our methodology.
  3. (3)
@@
    We validate our methodology with extensive experiments, demonstrating its potential to significantly enhance valuation of data, improve the configuration of models, and in this manner strengthen regulatory enforcement of global supply chains.
2. Related Work
@@
SIRA has been widely employed to discriminate geographic origin of various plant and animal-based products in global supply chains, such as garlic (Pianezze et al., 2019), Chinese tea (Liu et al., 2020), olive oil (Bontempo et al., 2019), cheese (Camin et al., 2004), and timber (Mortier et al., 2024; Truszkowski et al., 2025). By bringing data valuation methods to bear upon SIRA pipelines we aim to improve the verification of product provenance.
@@
Prior work in data valuation is typically seen in the context of explainable machine learning and enhancing model performance (Wu et al., 2024; Covert et al., 2024). Existing methods primarily rely on leave-one-out retraining and influence functions (Koh and Liang, 2017), Shapley values (Jia et al., 2019; Ghorbani and Zou, 2019; Wang and Jia, 2023), Least Cores (Yan and Procaccia, 2021), the Banzhaf value (Wang and Jia, 2022), Beta Shapley (Kwon and Zou, 2021), and reinforcement learning (Yoon et al., 2020). Furthermore, data valuation has been applied across various domains to enhance model development and interpretability, including health data (Pandl et al., 2021), medical imaging (Tang et al., 2021), and the Internet of Things (Shi and Duan, 2024). This paper is the first to formally apply data valuation techniques to SIRA.
3. Methods
3.1. Notation
@@
Let $\mathcal{X}$ be a set of locations where data is collected, where $x\in\mathcal{X}$ typically is specified by a longitude and latitude. Let $\mathcal{Y}$ be the set of measurements made over $\mathcal{X}$, here denoting stable isotope ratio values (e.g. $\delta^{13}C$, $\delta^{2}H$, $\delta^{15}N$, $\delta^{18}O$, $\delta^{34}S$), or trace element values (e.g. Si, Cu, S, Ba, Rb). We denote $f$ and $g$ as functions of interest (defined below). For evaluation purposes, we split our data into training and test datasets, where $D=\{z_{i}\}_{i=1}^{N}$ represents the training dataset with $N$ data points and $z_{i}=(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$. Similarly, $T=\{z_{i}\}_{i=1}^{M}$ denotes the test dataset with $M$ data points. We let $v(h,A,B)$ denote the performance of the model $h$ trained on a dataset $A$ and evaluated on the dataset $B$, where $v$ would return a numerical value, $v(h,A,B)\in\mathbb{R}$. In cases when the function and the test dataset are known, we drop the dependence in the notation to simply say $v(A)$.
3.2. Forward and Backward Models
@@
For the forward model, the task is to predict the stable isotope values for a given location, i.e., $f:\mathcal{X}\rightarrow\mathcal{Y}$. The motivation is to verify whether the characteristics of the location (denoting specified harvest origin) would align with general isotopic values associated with the specified location. Such models can be decision trees, random forests, or XGBoost to fit such functions. Recent works have proposed using Gaussian process regression models with high performance (Truszkowski et al., 2025; Mortier et al., 2024).
@@
In the backward case, one wants to identify the location given measured stable isotope values. We model this relation as $g:\mathcal{Y}\rightarrow\mathcal{X}$. The fitted model $g$ would help identify whether the declared harvest location of a species sample aligns with the predicted location from $g$. Similarly, here, one can use a range of machine learning models to fit this function. Moreover, Mortier et al. (2024) reversed the fitted Gaussian process regression models using the Bayes’ rule to predict locations from measured isotope ratios.
@@
Random Forest Models.
Random forests (RF) offer a robust and versatile approach for both forward and backward modeling in this context. As an ensemble learning method, it leverages the power of multiple decision trees to improve prediction accuracy and reduce overfitting. For the forward model $f:\mathcal{X}\rightarrow\mathcal{Y}$, an RF constructs $B$ decision trees $\{f_{t}\}_{t=1}^{B}$. Each tree $f_{t}$ is trained on a bootstrap sample $D_{t}$ of the training dataset $D$, where $D_{t}$ is obtained by randomly sampling $N$ data points from $D$ with replacement. Furthermore, during the construction of each tree $f_{t}$, specifically at each node split, a random subset of features $F\subseteq\mathcal{X}$ is considered from the full feature space $\mathcal{X}$. Let $M$ be the number of features randomly sampled at each split ($M<|\mathcal{X}|$). For a given location $x\in\mathcal{X}$, each tree $f_{t}$ provides a prediction $f_{t}(x)$. The final RF prediction $\hat{f}(x)$ is then obtained by aggregating these individual predictions, typically through averaging for regression tasks:
$\hat{f}(x)=\frac{1}{B}\sum_{t=1}^{B}f_{t}(x;D_{t},F),$
where $f_{t}(x;D_{t},F)$ denotes the prediction of the $t$-th tree trained on bootstrap dataset $D_{t}$ and using random feature subsets $F$ at each split. Similarly, for the backward model $g:\mathcal{Y}\rightarrow\mathcal{X}$, an RF comprised of $B$ trees $\{g_{t}\}_{t=1}^{B}$ can be trained. Each tree $g_{t}$ is trained on a bootstrap sample $D_{t}$ of $D$ to predict the location $x\in\mathcal{X}$ based on features $y\in\mathcal{Y}$, again using random feature subsets from $\mathcal{Y}$ at each split. The ensemble prediction for the backward model $\hat{g}(y)$ given a feature vector $y\in\mathcal{Y}$ is calculated as:
$\hat{g}(y)=\frac{1}{B}\sum_{t=1}^{B}g_{t}(y;D_{t},F^{\prime})$
Here, $g_{t}(y;D_{t},F^{\prime})$ represents the prediction of the $t$-th tree for the backward model, trained on bootstrap sample $D_{t}$ and using random feature subsets $F^{\prime}\subseteq\mathcal{Y}$ during tree construction. The performance $v(\hat{f},D,T)$ and $v(\hat{g},D,T)$ can then be evaluated for the forward and backward random forest models, respectively, using the provided metric $v$.
@@
Gaussian process regression (GPR) models, as explored by (Truszkowski et al., 2025; Mortier et al., 2024), offer a powerful approach for both forward and backward modeling settings. For the forward model $f:\mathcal{X}\rightarrow\mathcal{Y}$, GPR can be used to predict isotope ratios $y\in\mathcal{Y}$ at a given location $x\in\mathcal{X}$. We can construct “isoscapes” by fitting independent GP regression models to each feature in $\mathcal{Y}$. Considering our training dataset $D=\{z_{i}\}_{i=1}^{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$, for each feature $y_{j}\in\mathcal{Y}$, a GP model is trained to predict isotope values at a new location $x^{*}\in\mathcal{X}$. The predicted distribution for the feature $y_{j}$ at location $x^{*}$ is Gaussian, with a mean and variance given by:
$E[y_{j}|x^{*},\mathbf{X}]=\mu_{j}+\mathbf{k}^{(j)T}(\mathbf{K}^{(j)}+\sigma_{j}^{2}\mathbf{I})^{-1}(\mathbf{y}_{j}-\mu_{j}),$
$V(y_{j}|x^{*},\mathbf{X})=k^{(j)}(\mathbf{x^{*}},\mathbf{x^{*}})+\sigma_{j}^{2}-\mathbf{k}^{(j)T}(\mathbf{K}^{(j)}+\sigma_{j}^{2}\mathbf{I})^{-1}\mathbf{k}^{(j)T}.$
@@
$p(y^{*}|x^{*},D)=\prod_{j\in\mathcal{Y}}\frac{1}{\sqrt{2\pi V(y_{j}|x^{*},D)}}\exp\left(-\frac{(y_{j}^{*}-E[y_{j}|x^{*},D])^{2}}{2V(y_{j}|x^{*},D)}\right).$
@@
[…]
@@
3.3. Data Valuation
Data valuation is crucial for understanding the contribution of individual training samples to the performance of trained models for timber provenance. The Shapley value, introduced by (Shapley, 1953), offers a principled approach to quantify this contribution, identifying both highly informative and potentially detrimental data points. The Shapley value, $\phi_{i}$, for a data point $i$ is computed as the weighted average of its marginal contribution to model performance across all possible subsets of the training data:
@@ 3.3. Data Valuation
Here, $D$ is the full training set, $S$ is a subset excluding $i$, and $v(S)$ is the model performance (e.g., negative mean absolute error) when trained on subset $S$. High positive Shapley values indicate highly valuable data points that significantly improve performance, while low or negative values suggest redundancy or detrimental effects, possibly due to outliers, measurement errors, or model misspecification. The Shapley value is not an arbitrary metric; it is uniquely characterized by satisfying a set of desirable axioms, ensuring fairness and consistency in data valuation:
  1. (1)
@@
  4. (4)
    Dummy: if two data points $i$ and $j$ always have the same marginal contribution to every subset of $D$ then their Shapley value must be equal. $\phi_{i}=\phi_{j}$.
These axioms provide a strong theoretical justification for using the Shapley value. Furthermore, the Shapley value can be equivalently expressed as a sum over permutations of the dataset (Shapley, 1953):

### v3

@@ Abstract
Determining and verifying product provenance remains a critical challenge in global supply chains, particularly as geopolitical conflicts and shifting borders create new incentives for misrepresentation of commodities, such as hiding the origin of illegally harvested timber or stolen agricultural products. Stable Isotope Ratio Analysis (SIRA), combined with Gaussian process regression-based isoscapes, has emerged as a powerful tool for geographic origin verification. While these models are now actively deployed in operational settings supporting regulators, certification bodies, and companies, they remain constrained by data scarcity and suboptimal dataset selection. In this work, we introduce a novel deployed data valuation framework designed to enhance the selection and utilization of training data for machine learning models applied in SIRA. By quantifying the marginal utility of individual samples using Shapley values, our method guides strategic, cost-effective, and robust sampling campaigns within active monitoring programs. By prioritizing high-informative samples, our approach improves model robustness and predictive accuracy across diverse datasets and geographies. Our framework has been implemented and validated in a live provenance verification system currently used by enforcement agencies, demonstrating tangible, real-world impact. Through extensive experiments and deployment in a live provenance verification system, we show that this system significantly enhances provenance verification, mitigates fraudulent trade practices, and strengthens regulatory enforcement of global supply chains.
1 Introduction
@@
Global natural resource supply chains are opaque, especially since natural resources are often transformed from raw materials (e.g., timber) into finished consumer-facing products (e.g., furniture). These complex supply chains often involve multiple countries, with intermediate outputs being traded internationally and being used as inputs into further manufacturing processes. In addition to business-commerce decisions driving supply chain sourcing, the economics of natural resource trade are often closely linked with geopolitics. Verifying the true origin of products remains difficult and high-stakes, as geopolitical incentives and “don’t ask, don’t tell” sourcing cultures encourages misrepresentation of commodities, particularly for illegally harvested timber or stolen agricultural products.
The Scale of the Problem: Illegal Timber as a Case Study.
Illegal logging, in particular, is the most profitable natural resource crime, valued at US\$52 billion to US\$157 billion per year (May 2017). Illegally obtained timber accounts for 10–30% of the total global trade in timber products, and in regions such as Southeast Asia, Central Africa, and South America, it is estimated that 50–90% of timber is harvested illegally (May 2017). These figures underscore the urgent need for deployed, verifiable, and scientifically defensible provenance systems.
Forensic and Analytical Tools for Provenance Verification.
Product identification and provenance verification of traded natural resources have emerged as promising research areas, with various combinations of methods used depending on the resource sector and the granularity required for species identification and origin determination; for example, species and geographic harvest provenance for wood and forest products often requires multiple forensic tools and testing methods (woo 2021; Schmitz et al. 2020; Dormontt et al. 2015). For geographic origin verification in particular, Stable Isotope Ratio Analysis (SIRA), combined with Gaussian process regression-based isoscapes, has proven highly effective (Truszkowski et al. 2025; Mortier et al. 2024). SIRA leverages the natural variation in stable isotope ratios—measured via mass spectrometry—to determine the enrichment of non-radioactive elemental isotopes in a sample (Barrie and Prosser 1996), with enrichment patterns driven by environmental, atmospheric, soil, metabolic, and species-specific factors (Siegwolf et al. 2022; Wang et al. 2021; Vystavna, Matiatos, and Wassenaar 2021). It has been successfully used to trace the origin of timber, seafood, agricultural products, and fiber such as cotton (Truszkowski et al. 2025; Mortier et al. 2024; Watkinson et al. 2022; Cusa et al. 2022; Wang et al. 2020; Meier-Augenstein et al. 2014). However, operational deployment of these models faces challenges of data scarcity, suboptimal reference selection, and high sampling costs, motivating the need for systematic data valuation and prioritization. This in turn highlights the importance of optimizing and ‘valuing’ reference sample collection efforts (Gasson et al. 2021).
Operationalizing Provenance Verification: The WFID Platform.
Efforts such as World Forest ID (WFID) (worldforestid.org) have operationalized geographic origin models to address this problem, particularly in the timber trade. WFID identifies high-risk species and geographies (Norman 2023), collects chain-of-custody reference samples, and uses laboratory analysis (Fera Science 2025) to generate the chemical data needed for stable isotope modeling. These models are not only research prototypes—they are currently deployed and used in live compliance systems. The World Forest ID Evaluation Platform (World Forest ID 2024b) allows regulators, certification bodies, and companies to evaluate sourcing claims in real time.
Here, we introduce our deployed data valuation framework that enhances the selection and utilization of training data for machine learning models applied to SIRA. Field sample collections, central to any scientific traceability method, are logistically difficult and expensive due to remote locations, specialized equipment, and labor-intensive workflows. Our deployed system operationalizes Shapley-based data valuation to quantify the marginal utility of individual samples, enabling strategic, cost-effective, and robust sampling campaigns that directly improve model accuracy in active enforcement workflows.
Contributions:
  1. 1.
@@
    Application of machine learning data valuation to an operational provenance verification context. By prioritizing highly informative samples, our approach improves model robustness and predictive accuracy across diverse datasets and geographies.
  2. 2.
@@
    Deployment in live enforcement systems. We have deployed our optimized sampling approach in a provenance verification system used by European enforcement agencies to curb the trade of sanctioned Russian timber by proving that alternate claimed origins are non-viable. See coverage of our work in the New York Times (Nazaryan 2024)). Due to confidentiality reasons, we use a global dataset of Oak (Quercus spp.) reference samples to illustrate our methodology.
  3. 3.
@@
    Extensive empirical validation and field integration. We validate our framework with global Oak (Quercus spp.) reference datasets, demonstrating its potential to enhance data valuation, optimize model configuration, and strengthen the regulatory enforcement of global supply chains.
2 Related Work
@@
SIRA has been widely employed as a geographical discriminator for various plant and animal-based products in global supply chains, such as garlic (Pianezze et al. 2019), Chinese tea (Liu et al. 2020), olive oil (Bontempo et al. 2019), cheese (Camin et al. 2004), and timber (Mortier et al. 2024; Truszkowski et al. 2025). By bringing data valuation methods to bear upon SIRA pipelines we aim to improve the verification of product provenance.
@@
Prior work in data valuation is typically seen in the context of explainable machine learning and enhancing model performance (Wu, Zhu, and Li 2024; Covert et al. 2024). Existing methods primarily rely on leave-one-out retraining and influence functions (Koh and Liang 2017), Shapley values (Jia et al. 2019; Ghorbani and Zou 2019; Wang and Jia 2023), Least Cores (Yan and Procaccia 2021), the Banzhaf value (Wang and Jia 2022), Beta Shapley (Kwon and Zou 2021), and reinforcement learning (Yoon, Arik, and Pfister 2020). Furthermore, data valuation has been applied across various domains to enhance model development and interpretability, including health data (Pandl et al. 2021), medical imaging (Tang et al. 2021), and the Internet of Things (Shi and Duan 2024). This paper is the first to formally apply data valuation techniques to SIRA.
3 Methods
@@
Let $\mathcal{X}$ be a set of locations where data is collected; $x\in\mathcal{X}$ typically is specified by a longitude and latitude. Let $\mathcal{Y}$ be the set of measurements made over $\mathcal{X}$, here denoting stable isotope ratio values (e.g. $\delta^{13}C$, $\delta^{2}H$, $\delta^{15}N$, $\delta^{18}O$, $\delta^{34}S$), or trace element values (e.g. Si, Cu, S, Ba, Rb). We denote $f$ and $g$ as functions of interest (defined below). For evaluation purposes, we split our data into training and test datasets, where $D=\{z_{i}\}_{i=1}^{N}$ represents the training dataset with $N$ data points and $z_{i}=(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$. Similarly, $T=\{z_{i}\}_{i=1}^{M}$ denotes the test dataset with $M$ data points. We let $v(h,A,B)$ denote the performance of the model $h$ trained on a dataset $A$ and evaluated on the dataset $B$, where $v$ would return a numerical value, $v(h,A,B)\in\mathbb{R}$. In cases when the function and the test dataset are known, we drop the dependence in the notation to simply say $v(A)$.
3.1 Forward and Backward Models
@@
For the forward model, the task is to predict the stable isotope values for a given location, i.e., $f:\mathcal{X}\rightarrow\mathcal{Y}$. The motivation is to verify whether the characteristics of the location (denoting specified harvest origin) would align with general isotopic values associated with the specified location. Such models can be based on approahes like decision trees, random forests, or XGBoost. Recent works have proposed using Gaussian process regression models with high performance (Truszkowski et al. 2025; Mortier et al. 2024).
@@
In the backward case, we aim to identify the location given measured stable isotope values. We model this relation as $g:\mathcal{Y}\rightarrow\mathcal{X}$. The fitted model $g$ would help identify whether the declared harvest location of a species sample aligns with the predicted location from $g$. Similarly, here, one can use a range of machine learning models to fit this function. For instance, Mortier et al. (2024) reversed a fitted Gaussian process regression model using the Bayes’ rule to predict locations from measured isotope ratios.
@@
Gaussian process regression (GPR) models, as explored by (Truszkowski et al. 2025; Mortier et al. 2024), offer a powerful approach for both forward and backward modeling settings. For the forward model $f:\mathcal{X}\rightarrow\mathcal{Y}$, GPR can be used to predict isotope ratios $y\in\mathcal{Y}$ at a given location $x\in\mathcal{X}$. We can construct “isoscapes” by fitting independent GP regression models to each feature in $\mathcal{Y}$. Considering our training dataset $D=\{z_{i}\}_{i=1}^{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$, for each feature $y_{j}\in\mathcal{Y}$, a GP model is trained to predict isotope values at a new location $x^{*}\in\mathcal{X}$. The predicted distribution for the feature $y_{j}$ at location $x^{*}$ is Gaussian, with a mean and variance given by:
$E[y_{j}|x^{*},\mathbf{X}]=$ $\mu_{j}+\mathbf{k}^{(j)T}(\mathbf{K}^{(j)}+\sigma_{j}^{2}\mathbf{I})^{-1}(\mathbf{y}_{j}-\mu_{j})$
$V(y_{j}|x^{*},\mathbf{X})=\;$ $k^{(j)}(\mathbf{x^{*}},\mathbf{x^{*}})+\sigma_{j}^{2}$
$-\mathbf{k}^{(j)T}(\mathbf{K}^{(j)}+\sigma_{j}^{2}\mathbf{I})^{-1}\mathbf{k}^{(j)}$
@@
$p(y^{*}\mid x^{*},D)$ $=\prod_{j\in\mathcal{Y}}\frac{1}{\sqrt{2\pi V(y_{j}\mid x^{*},D)}}$
$\quad\times\exp\left(-\frac{(y_{j}^{*}-E[y_{j}\mid x^{*},D])^{2}}{2V(y_{j}\mid x^{*},D)}\right)$
@@
$\text{RMSE}=\sqrt{\int_{\mathbf{x}\in A}\big(d(\mathbf{x}_{t},\mathbf{x})\big)^{2}\,p(\mathbf{x}|\mathbf{y}^{*},D)\,d\mathbf{x}},$
@@
3.2 Data Valuation
The Shapley value, introduced by (Shapley 1953), offers a principled approach to quantify data value, identifying both highly informative and potentially detrimental data points. The Shapley value $\phi_{i}$ for a data point $i$ is computed as the weighted average of its marginal contribution to model performance across all possible subsets of the training data:
@@ 3.2 Data Valuation
Here, $D$ is the full training set, $S$ is a subset excluding $i$, and $v(S)$ is the model performance (e.g., negative mean absolute error) when trained on subset $S$. High positive Shapley values indicate highly valuable data points that significantly improve performance, while low or negative values suggest redundancy or detrimental effects, possibly due to outliers, measurement errors, or model misspecification. The Shapley value is not an arbitrary metric; it is uniquely characterized as that satisfying a set of desirable axioms, ensuring fairness and consistency in data valuation:
  1. 1.
@@
  4. 4.
    Dummy: if two data points $i$ and $j$ always have the same marginal contribution to every subset of $D$ then their shapely value must be equal. $\phi_{i}=\phi_{j}$.
These axioms provide a strong theoretical justification for using the Shapley value. Furthermore, the Shapley value can be equivalently expressed as a sum over permutations of the dataset (Shapley 1953):